40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 13:22:13 2011, Leo, Update, SUS, Re: DRMI prep : suspension diagnostic 10x
    Reply  Tue May 3 15:04:20 2011, kiwamu, Update, SUS, Re: DRMI prep : suspension diagnostic 
    Reply  Wed May 4 13:57:04 2011, kiwamu, Update, SUS, mechanical resonances updated 
Message ID: 4613     Entry time: Tue May 3 15:04:20 2011     In reply to: 4611
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: Re: DRMI prep : suspension diagnostic 

Jenne went through all the suspension racks and pushed all the connectors.

After pushing them, we had a quick look at those spectra and found no funny noise spectrum except for C1:PRM-SENSOR_UL.

We then checked connection around the SCSI cables and eventually found the connection between ADC_card_0 and a SCSI was loose.

We put short standoffs on the ADC card so that the screws from the SCSI can nicely reach to the ADC card. Now everything looks fine.

SUS diagnostic is quite useful !

Quote from #4611

Notice that the C1:SUS-ITMX_SENSOR_UL and C1:SUS-MC3_SENSOR_UL spectra fall as 1/f.  Jenne suggested that this might indicate that there is a loose electrical connection.

ELOG V3.1.3-